# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 1. Introduction
The natural embedding of $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})$ into $\mathsf{D}\mathsf{(Cond(Ab))}$ is not full. In particular, the left hand side of the equation
$\mathrm{RHom}_{\mathsf{Pro}(\mathsf{D}^{\geq 0}(\mathsf{Ab}))}\left(\text{``}\prod_{\omega}\text{''}\,\bigoplus_{\omega_{1}}\mathbb{Z},\mathbb{Z}\right)\not\cong\mathrm{RHom}_{\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}}\left(\prod_{\omega}\bigoplus_{\omega_{1}}\mathbb{Z},\mathbb{Z}\right),$
@@
If, however, one restricts attention to countable abelian groups and operations upon them, then the main pathologies (so to speak) of Theorem A are avoidable. We show the following in Section 2.4; the $\mathbf{A}[H]$ appearing in its premise is the natural generalization of the inverse system $\mathbf{A}=\mathbf{A}[\mathbb{Z}]$ tracing back, via a substantial set theoretic literature, to [54].
@@
Let us underscore before proceeding any further that the forms of Theorems A, B, and D, as well as their conversions to questions about derived limits, are all essentially due to Clausen and Scholze; in each case, our own contributions to the theorems consist principally in the derived limit analyses (exemplified by our Theorems 2.10 and 4.10) which complete their proofs.
@@
  1. (1)
    For every field $K$ and $S\in\mathsf{ED}$, the constant sheaf $\mathcal{K}$ on $S$ is of injective dimension at most one.
  2. (2)
@@
    If $S,T\in\mathsf{ED}$ are each Čech-Stone compactifications of discrete sets of cardinality at least $\aleph_{\omega}$, then the injective dimension of the constant sheaf $\mathcal{K}$ on $S\times T$ is infinite.
  3. (3)
@@
    There exist compact projective condensed anima $S$ and $T$ whose product is not compact.
@@
Let us turn now in greater detail to the paper’s organization and the varying demands that it places upon its readers. Section 2 divides into five subsections; of these, only 2.1, which provides a brief review of derived and condensed settings, and 2.3, recording the fundamentals of the groups $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}$, play a significant role in subsequent sections. Subsections 2.2 and the more down-to-earth 2.4, containing main portions of the arguments of Theorems A and B, respectively, provide motivation and context for the computations of 2.3, while 2.5 records a more set theoretically involved refinement of the latter. Put differently, in Subsections 2.1, 2.3, and 2.4, we presume only a basic familiarity with homological algebra ([83, Ch. 1–3], say) and set theory, while in Subsections 2.2 and 2.5 (both of which may be skimmed in a first reading), we assume a little more familiarity, respectively, with each.
@@
Section 3 and 4 may be read independently of one another; the first of these divides into two subsections, neither of which presumes any further background of the reader. Its first subsection explores the implications of the results of Subsection 2.3 along the lines discussed before Theorem C above, which it also proves. A more optional second subsection contextualizes these implications within a broader discussion of strong homology.
@@
Next, after a brief introduction, the two subsections of Section 4 contain our proof of Theorem D, with the first containing the proofs of clauses (1) and (2) and the second containing that of clause (3). These portions of the paper require a bit more of the reader. Subsection 4.1 involves sheaf theory at roughly the level of [45, Ch. 2]. Subsection 4.2, like any discussion of condensed anima, then necessitates an at least occasionally $\infty$-category theoretic vocabulary; relevant references are provided along the way. The paper then concludes with a list of open questions; among these, those amounting essentially to problems in infinitary combinatorics predominate.
@@
We hope by this outline to have suggested the kinds of math which our text engages; its more fundamental roots in the more contemporary field of condensed mathematics, though, merit a few further words of comment. Core references for condensed mathematics are the Scholze and Clausen–Scholze lecture notes [74, 73, 23] and the online Clausen–Scholze lecture series [24] and [25]; multiple theses [58, 3, 53] and course notes [60, 72, 39] usefully supplement this material. In so rapidly expanding a field, however, reference lists take on a markedly provisional character. Contemporaneous with the development of condensed mathematics has been that of Barwick and Haine’s closely related pyknotic mathematics; among the fundamental references here are [10] and the MSRI lecture series [9], although our comments above on references proliferating faster than we can confidently track continue to apply. Below, we have uniformly framed our results in condensed terms, both for simplicity and for the essentially contingent reason that our first encounter with this world of ideas was by way of an email from Dustin Clausen and Peter Scholze; these results’ translations to the pyknotic framework are, in general, straightforward. For that initial contact and for much rich and generous conversation thereafter, we wish to underscore our thanks to Clausen and Scholze. For further discussion and clarifications of the material of Sections 2.2 and 4.2, we are grateful to Lucas Mann as well.
@@
Lastly, a word on some of our most fundamental notations. We write $P(X)$ for the power set, and $|X|$ for the cardinality, of a set $X$. If $X$ is partially ordered then we write $\mathrm{cf}(X)$ for the least cardinality of a cofinal subset of $X$, while if $\kappa$ is a cardinal then $\mathrm{Cof}(\kappa)$ denotes the class of ordinals $\xi$ for which $\mathrm{cf}(\xi)=\kappa$. If $X$ is well-ordered then $\mathrm{otp}(X)$ denotes the unique ordinal which is order-isomorphic to $X$. We write ${{}^{\kappa}}X$ for the collection of functions from $\kappa$ to $X$, although we will at times also exponentiate on the right. Relatedly, $[X]^{n}$ and $[X]^{<\omega}$ denote, respectively, the collections of size-$n$ subsets, and of finite subsets, of $X$; when $X$ is linearly ordered, we will identify these with ordered tuples in the natural way. If $\mathcal{C}$ is a category, then we write $X\in\mathcal{C}$ to indicate that $X$ is an object of $\mathcal{C}$.
@@ 2.1. A brief survey of the relevant categories
[…]
is well-defined on $\mathsf{Pro}(\mathcal{C})$. In particular (as a closer inspection of the morphisms of $\mathsf{Pro}(\mathcal{C})$ then implies), $\lim_{i\in J}C_{i}\cong\lim_{i\in I}C_{i}$ whenever $J$ is a cofinal subset of $I$. For more on the subject of pro-categories (often via their dual notion $\mathsf{Ind}(\mathcal{C}^{\mathrm{op}})=(\mathsf{Pro}(\mathcal{C}))^{\mathrm{op}}$, which plays, in turn, a significant conceptual role in Section 4 below), see [2, I.8], [1, Appendix], [48, VI], [56, I.1], or [49, §6].
@@
${\mathsf{K}^{\star}(\mathcal{A})}$${\mathsf{D}^{\star}(\mathcal{A})}$${\mathsf{K}^{\star}(\mathcal{B})}$${\mathsf{D}^{\star}(\mathcal{B})}$$\scriptstyle{\mathsf{K}(F)}$$\scriptstyle{\ell_{\mathcal{A}}}$$\scriptstyle{\ell_{\mathcal{B}}}$
@@
witnesses the failure of condition (4) above. The choice of $\mathbb{R}$ here is, of course, essentially arbitrary; the same issue arises for any fixed algebraic object admitting multiple topologies, some finer than others. These examples epitomize a tension between topological data and any practice of algebra partaking of the constructions of the previous paragraph; they are standard within the condensed literature and cue one of its main motivating questions: How to do algebra when rings/modules/groups carry a topology? [74, p. 6]. We turn now, with this as background, to a brief introductory overview of the condensed framework. This subsection’s remainder is entirely drawn from [74, I–II].
@@
where $\mathrm{Cont}(S,X)$ denotes the set of continuous maps from $S$ to $X$, is a condensed set, as the reader is encouraged to verify. In fact, the maps $X\mapsto\underline{X}$ define a fully faithful embedding of the category of compactly generated weakly Hausdorff topological spaces into the category $\mathsf{Cond}(\mathsf{Set})$; this is the first fundamental fact about the theory ([74, Prop. 1.7]). For the second, define the category $\mathsf{Cond}(\mathsf{Ab})$ either by replacing $\mathsf{Set}$ with $\mathsf{Ab}$ in Definition 2.1 or by restricting to the abelian group objects of $\mathsf{Cond}(\mathsf{Set})$. We then have [74, Thm. 1.10]:
@@
[…]
@@
A few remarks are immediately in order. First: as, in general, it’s only at the $\infty$-category theoretic level that the $\mathsf{Pro}(-)$ operation preserves derived structure (cf. [49, §15.4]), it’s at this level that Question 2.3 should ultimately be posed. We won’t belabor this point, since it only momentarily affects our reformulation of the question below; more generally, as our introduction suggested, a relaxed or “naive” reading relationship (cf. [22, Ch. I.1]; the analogy, of course, is with naive set theory) to the $\infty$-category theoretic manipulations of both this subsection and Section 4.2 is broadly compatible with this paper’s main aims. This is fortunate, since any rigorous review of the subject of $\infty$-categories is beyond the scope of this paper. Lurie’s [50] or its more recent online reworking [52] are standard one-stop references; for the subject of $\infty$-derived categories, see [51, §1.3]. For a highly economical review of much of that material, see Appendix A of [43], which we cite more particularly in Section 4 below. In any case, a main point in the present context is that, when interpreted at the $\infty$-derived level, the $\mathrm{Rlim}$ of (3) is simply a limit; thus at this level, the functor (3) is plainly limit-preserving, and Question 2.3 is, modulo size issues, one of whether $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})^{\mathtt{b}}$ is a reflective subcategory of $\mathsf{D}\mathsf{(Cond(Ab))}$ [50, §5.2.7]. [^1] The issues in question are that the constituent categories aren’t presentable, so that the standard lemma [50, Cor. 5.5.2.9] completing the characterization doesn’t apply. We will expand on the question’s significance in this subsection’s conclusion below.
@@
see [49, 2.6.2]. Here $\mathbb{Z}[-n]$ denotes the cochain complex valued $\mathbb{Z}$ in the degree $n$ and 00 elsewhere and the natural map $\bigoplus_{\mathbb{N}}\mathbb{Z}[-n]\to\prod_{\mathbb{N}}\mathbb{Z}[-n]$ is easily seen to be a quasi-isomorphism, with the consequence that
@@
In each of the above cases the concluding $\mathrm{Hom}$ terms are $\mathrm{H}^{n}(E)$, hence if the cohomology of $E$ is nonvanishing in unboundedly many degrees, the two leftmost terms are non-isomorphic, and this instance of the functor (3) isn’t full. Restricting attention to the full subcategory of $\mathsf{Pro(D(Ab))}$ spanned by inverse systems of uniformly bounded complexes, though, averts this issue, and it is this category which we have denoted above by $\mathsf{Pro(D(Ab))}^{\mathtt{b}}$. So posed, Question 2.3 reduces to a more computationally concrete and interesting question, by steps which we now describe.
First, note that one may without loss of generality restrict to complexes bounded from below by 00, and hence to the derived category $\mathsf{D}^{\geq 0}$; observe next that by equation 1, our question is then that of whether for all inverse systems of uniformly bounded cochain complexes $(C_{i})_{i\in I}$ and $(E_{j})_{j\in J}$ in $\mathsf{D}^{\geq 0}(\mathsf{Ab})$ the natural map
@@
is an isomorphism for every bounded $E$ in $\mathsf{D}^{\geq 0}(\mathsf{Ab})$ (note that by the example above, this is in general false if $E$ is unbounded). By passing the $\mathop{colim}\displaylimits$ back across the parentheses and recalling that $\mathrm{RHom}$ commutes with all finite limits and colimits, we may by [50, 4.4.2.7] further reduce our question to that of whether
@@
for all $(C_{i})_{i\in I}$ and $E$ as above. Standard dévissage arguments (cf. [49, Prop. 15.4.2]) using the truncation functors $\tau^{\geq n}$ and fiber sequences $\tau^{<n}E\to E\to\tau^{\geq n}E$ allow us to further reduce to the case in which $E$ is concentrated in a single degree; similarly for the first coordinate. Working with $\mathrm{RHom}$ allows us to further assume, without loss of generality, that that degree in both cases is zero. We have thus reduced our question to the following.
@@
Recall that $[J]^{<\omega}$ denotes the collection of finite subsets of $J$. For any function $f:I\to[J]^{<\omega}$, let $X(f)=\{(i,j)\in I\times J\mid j\in f(i)\}$, and let ([${{}^{I}}([J]^{<\omega})$ denote the ordering of functions from $I$ to $[J]^{<\omega}$ given by $f\leq g$ if and only if $X(f)\subseteq X(g)$. Now observe that
@@
(cf. [68, Prop. 3.6.3]). Observe now that, since $\prod_{i\in I}\underline{\mathbb{Z}^{f(i)}}$ is projective in the subcategory $\mathsf{Solid}$ of $\mathsf{Cond(Ab)}$, and that the derived category of the former fully and faithfully embeds in that of the latter ([74, Thm. 5.8]), the interior $\mathrm{RHom}$ above concentrates in degree zero, and may be computed in that degree to equal $\bigoplus_{i\in I}\bigoplus_{f(i)}H=\bigoplus_{X(f)}H$. This last point follows essentially from [36, Prop. 97.7 or Cor. 94.5]. We arrive in this way to the derived limits of the inverse systems $\mathbf{A}_{I,J}[H]$, the focus of the following subsection. As will soon grow clear, these limits are substantially more complicated than the left-hand side expressions of line 4; as such, they exemplify the computational import of Question 2.3: it is asking whether, on a significant subcategory of $\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}$, the $\mathrm{RHom}_{\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}}$ functor admits a more transparent or manageable description. That subcategory $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})^{\mathtt{b}}$ is significant not only for formalizing questions about limits, but for its centrality (hinted at above) to the intermediate subcategory $\mathsf{D}^{\geq 0}(\mathsf{Solid})$, one of the most critical in condensed mathematics (see [74, §5, 6], [73, §2]). In essence, the above conversion was via the case of finite $J$, and the rather nontrivial computation of that case is a foundational result in the theory of solid abelian groups. Question 2.3 may be read as one of how that computation extends.

### v2

@@ 1. Introduction
The natural functor sending $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})$ to $\mathsf{D}\mathsf{(Cond(Ab))}$ is not full. In particular, the left hand side of the equation
$\mathrm{RHom}_{\mathsf{Pro}(\mathsf{D}^{\geq 0}(\mathsf{Ab}))}\left(\text{``}\prod_{\omega}\text{''}\,\bigoplus_{\omega_{1}}\mathbb{Z},\mathbb{Z}\right)\not\cong\mathrm{RHom}_{\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}}\left(\prod_{\omega}\bigoplus_{\omega_{1}}\mathbb{Z},\mathbb{Z}\right)$
@@
If, however, one restricts to countable abelian groups and operations upon them, then the main pathologies (so to speak) of Theorem A are avoidable. We show the following in Section 2.4; the $\mathbf{A}[H]$ appearing in its premise is the natural generalization of the inverse system $\mathbf{A}=\mathbf{A}[\mathbb{Z}]$ tracing back, via a substantial set theoretic literature, to [57].
@@
Let us underscore before proceeding any further that the forms of Theorems A, B, and D, as well as their conversions to questions about derived limits, are all essentially due to Clausen and Scholze; in each case, our own contributions to the theorems consist principally in the derived limit analyses (exemplified by our Theorems 2.10 and 4.11) which complete their proofs.
@@
For every finite field $K$ and $S\in\mathsf{ED}$, the constant sheaf $\mathcal{K}$ on $S$ is injective. If $S,T\in\mathsf{ED}$ are each Čech-Stone compactifications of discrete sets of cardinality at least $\aleph_{\omega}$, then for any field $K$ the injective dimension of the constant sheaf $\mathcal{K}$ on $S\times T$ is infinite.
@@
In consequence, there exist compact projective condensed anima $S$ and $T$ whose product is not compact.
@@
Let us turn now in greater detail to the paper’s organization and the varying demands that it places upon its readers. Section 2 divides into five subsections; of these, only 2.1, which provides a brief review of derived and condensed settings, and 2.3, recording the fundamentals of the groups $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}$, play a significant role in subsequent sections. Subsections 2.2 and the more down-to-earth 2.4, containing main portions of the arguments of Theorems A and B, respectively, provide motivation and context for the computations of 2.3, while 2.5 records a more set theoretically involved refinement of the latter. Put differently, in Subsections 2.1, 2.3, and 2.4, we presume only a basic familiarity with homological algebra ([86, Ch. 1–3], say) and set theory, while in Subsections 2.2 and 2.5 (both of which may be skimmed in a first reading), we assume a little more familiarity, respectively, with each.
@@
Sections 3 and 4 may be read independently of one another; the first of these divides into two subsections, neither of which presumes any further background of the reader. Its first subsection explores the implications of the results of Subsection 2.3 along the lines discussed before Theorem C above, which it also proves. A more optional second subsection contextualizes these implications within a broader discussion of strong homology.
@@
Next, after a brief introduction, the two subsections of Section 4 contain our proof of Theorem D, with the first containing the proofs of its first paragraph and the second containing that of its second. These portions of the paper require a bit more of the reader. Subsection 4.1 involves sheaf theory at roughly the level of [48, Ch. 2]. Subsection 4.2, like any discussion of condensed anima, then necessitates an at least occasionally $\infty$-category theoretic vocabulary; relevant references are provided along the way. The paper then concludes with a list of open questions; among these, those amounting essentially to problems in infinitary combinatorics predominate.
@@
We hope by this outline to have suggested the kinds of math which our text engages; its more fundamental roots in the more contemporary field of condensed mathematics, though, merit a few further words of comment. Core references for condensed mathematics are the Scholze and Clausen–Scholze lecture notes [77, 76, 25] and the online Clausen–Scholze lecture series [26] and [27]; multiple theses [61, 3, 56] and course notes [63, 75, 42] usefully supplement this material. In so rapidly expanding a field, however, reference lists take on a markedly provisional character. Contemporaneous with the development of condensed mathematics has been that of Barwick and Haine’s closely related pyknotic mathematics; among the fundamental references here are [11] and the MSRI lecture series [10], although our comments above on references proliferating faster than we can confidently track them continue to apply. Below, we have uniformly framed our results in condensed terms, both for simplicity and for the essentially contingent reason that our first encounter with this world of ideas was by way of an email from Dustin Clausen and Peter Scholze; these results’ translations to the pyknotic framework are, in general, straightforward. For that initial contact and for much rich and generous conversation thereafter, we wish to underscore our thanks to Clausen and Scholze. For further discussion and clarifications of the material of Sections 2.2 and 4.2, we are grateful to Lucas Mann as well.
@@
Lastly, a word on some of our most fundamental notations. We write $P(X)$ for the power set, and $|X|$ for the cardinality, of a set $X$. If $X$ is partially ordered then we write $\mathrm{cf}(X)$ for the least cardinality of a cofinal subset of $X$, while if $\kappa$ is a cardinal then $\mathrm{Cof}(\kappa)$ denotes the class of ordinals $\xi$ for which $\mathrm{cf}(\xi)=\kappa$. If $\beta$ is an ordinal, then $\lim(\beta)$ will denote the set of all limit ordinals less than $\beta$. We note that we also use $\lim$ to denote the inverse limit functor; the correct interpretation will always be clear from context. If $X$ is well-ordered then $\mathrm{otp}(X)$ denotes the unique ordinal which is order-isomorphic to $X$. We write ${{}^{\kappa}}X$ for the collection of functions from $\kappa$ to $X$, although we will at times also exponentiate on the right. Relatedly, $X]^{n}$ and $X]^{<\omega}$ denote, respectively, the collections of size-$n$ subsets, and of finite subsets, of $X$; when $X$ is linearly ordered, we will identify these with ordered tuples in the natural way. If $\mathcal{C}$ is a category, then we write $X\in\mathcal{C}$ to indicate that $X$ is an object of $\mathcal{C}$.
@@ 2.1. A brief survey of the relevant categories
$\textnormal{``}\lim_{i\in I}\!\textnormal{''}\,C_{i}\,\mapsto\,\lim_{i\in I}C_{i}$
is well-defined on $\mathsf{Pro}(\mathcal{C})$. In particular (as a closer inspection of the morphisms of $\mathsf{Pro}(\mathcal{C})$ then implies), $\lim_{i\in J}C_{i}\cong\lim_{i\in I}C_{i}$ whenever $J$ is a cofinal subset of $I$. For more on the subject of pro-categories (often via their dual notion $\mathsf{Ind}(\mathcal{C}^{\mathrm{op}})=(\mathsf{Pro}(\mathcal{C}))^{\mathrm{op}}$, which plays, in turn, a significant conceptual role in Section 4 below), see [2, I.8], [1, Appendix], [51, VI], [59, I.1], or [52, §6].
@@
$\mathsf{K}^{\star}(\mathcal{A})$$\mathsf{D}^{\star}(\mathcal{A})$$\mathsf{K}^{\star}(\mathcal{B})$$\mathsf{D}^{\star}(\mathcal{B})$$\scriptstyle{\mathsf{K}(F)}$$\scriptstyle{\ell_{\mathcal{A}}}$$\scriptstyle{G}\scriptstyle{\ell_{\mathcal{B}}}$
@@
witnesses the failure of condition (4) above. The choice of $\mathbb{R}$ here is, of course, essentially arbitrary; the same issue arises for any fixed algebraic object admitting multiple topologies, some finer than others. These examples epitomize a tension between topological data and any practice of algebra partaking of the constructions of the previous paragraph; they are standard within the condensed literature and cue one of its main motivating questions: How to do algebra when rings/modules/groups carry a topology? [77, p. 6]. We turn now, with this as background, to a brief introductory overview of the condensed framework. This subsection’s remainder is entirely drawn from [77, I–II].
@@
where $\mathrm{Cont}(S,X)$ denotes the set of continuous maps from $S$ to $X$, is a condensed set, as the reader is encouraged to verify. In fact, the maps $X\mapsto\underline{X}$ define a fully faithful embedding of the category of compactly generated weakly Hausdorff topological spaces into the category $\mathsf{Cond}(\mathsf{Set})$; this is the first fundamental fact about the theory ([77, Prop. 1.7]). For the second, define the category $\mathsf{Cond}(\mathsf{Ab})$ either by replacing $\mathsf{Set}$ with $\mathsf{Ab}$ in Definition 2.1 or by restricting to the abelian group objects of $\mathsf{Cond}(\mathsf{Set})$. We then have [77, Thm. 1.10]:
@@
(3) $\mathrm{``lim}\textnormal{''}\,C_{i}\mapsto\mathrm{Rlim}\,\underline{C}_{i}$
@@
A few remarks are immediately in order. First: as, in general, it’s only at the $\infty$-category theoretic level that the $\mathsf{Pro}(-)$ operation preserves derived structure (cf. [52, §15.4]), it’s at this level that Question 2.3 should ultimately be posed. We won’t belabor this point, since it only momentarily affects our reformulation of the question below; more generally, as our introduction suggested, a relaxed or “naive” reading relationship (cf. [24, Ch. I.1]; the analogy, of course, is with naive set theory) to the $\infty$-category theoretic manipulations of both this subsection and Section 4.2 is broadly compatible with this paper’s main aims. This is fortunate, since any rigorous review of the subject of $\infty$-categories is beyond the scope of this paper. Lurie’s [53] or its more recent online reworking [55] are standard one-stop references; for the subject of $\infty$-derived categories, see [54, §1.3]. For a highly economical review of much of that material, see Appendix A of [46], which we cite more particularly in Section 4 below. In any case, a main point in the present context is that, when interpreted at the $\infty$-derived level, the $\mathrm{Rlim}$ of (3) is simply a limit; thus at this level, the functor (3) is plainly limit-preserving, and Question 2.3 is, modulo size issues, one of whether $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})^{\mathtt{b}}$ is a reflective subcategory of $\mathsf{D}\mathsf{(Cond(Ab))}$ [53, §5.2.7]. [^1] The issues in question are that the constituent categories aren’t presentable, so that the standard lemma [53, Cor. 5.5.2.9] completing the characterization doesn’t apply. We will expand on the question’s significance in this subsection’s conclusion below.
@@
see [52, 2.6.2]. Here $\mathbb{Z}[-n]$ denotes the cochain complex valued $\mathbb{Z}$ in the degree $n$ and 0 elsewhere and the natural map $\bigoplus_{\mathbb{N}}\mathbb{Z}[-n]\to\prod_{\mathbb{N}}\mathbb{Z}[-n]$ is easily seen to be a quasi-isomorphism, with the consequence that
@@
In each of the above cases the concluding $\mathrm{Hom}$ terms are $\mathrm{H}^{n}(E)$, hence if the cohomology of $E$ is nonvanishing in unboundedly many degrees, the two leftmost terms above are non-isomorphic, and this instance of the functor (3) isn’t full. Restricting attention to the full subcategory of $\mathsf{Pro(D(Ab))}$ spanned by inverse systems of uniformly bounded complexes, though, averts this issue, and it is this category which we have denoted above by $\mathsf{Pro(D(Ab))}^{\mathtt{b}}$. So posed, Question 2.3 reduces to a more computationally concrete and interesting question, by steps which we now describe.
First, note that one may without loss of generality restrict to complexes bounded from below by 0, and hence to the derived category $\mathsf{D}^{\geq 0}$; observe next that by equation 1, our question is then that of whether for all inverse systems of uniformly bounded cochain complexes $(C_{i})_{i\in I}$ and $(E_{j})_{j\in J}$ in $\mathsf{D}^{\geq 0}(\mathsf{Ab})$ the natural map
@@
is an isomorphism for every bounded $E$ in $\mathsf{D}^{\geq 0}(\mathsf{Ab})$ (note that by the example above, this is in general false if $E$ is unbounded). By passing the $\mathop{colim}\displaylimits$ back across the parentheses and recalling that $\mathrm{RHom}$ commutes with all finite limits and colimits, we may by [53, 4.4.2.7] further reduce our question to that of whether
@@
for all $(C_{i})_{i\in I}$ and $E$ as above. Standard dévissage arguments (cf. [52, Prop. 15.4.2]) using the truncation functors $\tau^{\geq n}$ and fiber sequences $\tau^{<n}E\to E\to\tau^{\geq n}E$ allow us to further reduce to the case in which $E$ is concentrated in a single degree; similarly for the first coordinate. Working with $\mathrm{RHom}$ allows us to further assume, without loss of generality, that that degree in both cases is zero. We have thus reduced our question to the following.
@@
[…]
@@
(cf. [71, Prop. 3.6.3]). Observe now that, since $\prod_{i\in I}\underline{\mathbb{Z}^{f(i)}}$ is projective in the subcategory $\mathsf{Solid}$ of $\mathsf{Cond(Ab)}$, and that the derived category of the former fully and faithfully embeds in that of the latter ([77, Thm. 5.8]), the interior $\mathrm{RHom}$ above concentrates in degree zero, and may be computed in that degree to equal $\bigoplus_{i\in I}\bigoplus_{f(i)}H=\bigoplus_{X(f)}H$. This last point follows essentially from [39, Prop. 97.7 or Cor. 94.5]. We arrive in this way to the derived limits of the inverse systems $\mathbf{A}_{I,J}[H]$, the focus of the following subsection. As will soon grow clear, these limits are substantially more complicated than the left-hand side expressions of line 4; as such, they exemplify the computational import of Question 2.3: it is asking whether, on a significant subcategory of $\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}$, the $\mathrm{RHom}_{\mathsf{D}^{\geq 0}\mathsf{(Cond(Ab))}}$ functor admits a more transparent or manageable description. That subcategory $\mathsf{Pro}(\mathsf{D}\mathsf{(Ab)})^{\mathtt{b}}$ is significant not only for formalizing questions about limits, but for its centrality (hinted at above) to the intermediate subcategory $\mathsf{D}^{\geq 0}(\mathsf{Solid})$, one of the most critical in condensed mathematics (see [77, §5, 6], [76, §2]). In essence, the above conversion was via the case of finite $J$, and the rather nontrivial computation of that case is a foundational result in the theory of solid abelian groups. Question 2.3 may be read as one of how that computation extends.